# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ Introduction
Let $k,N\geq 1$ be integers. We denote by $\mathcal{S}_{k}(\Gamma_{1}(N))$ the space of cuspidal holomorphic modular forms with level $\Gamma_{1}(N)$. Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a cuspidal newform with character $\varepsilon_{f}$, and $L$ be a number field containing its coefficients. Given a maximal ideal $\mathfrak{p}\subset\mathcal{O}_{L}$, we can attach to $f$ a free $\mathcal{O}_{L_{\mathfrak{p}}}$-module $T_{f,\mathfrak{p}}$ of rank two endowed with a continuous linear action of $G_{{\mathbb{Q}}}:=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})$. Works of Momose [17] and Ribet [21] show that, when $f$ has no complex multiplication and $k\geq 2$, the image of $G_{{\mathbb{Q}}}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\simeq\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$ contains a conjugate of $\operatorname{SL}_{2}({\mathbb{Z}}_{p})$ for all but finitely many $\mathfrak{p}$. More recently, Loeffler [14] proved an adelic open image theorem for such newforms, the analog of that of Serre [23] for elliptic curves without complex multiplication.
@@
then, for all but finitely many $\mathfrak{p}$, there exists a $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}:=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}}(\mu_{p^{\infty}}))$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one. Loeffler also discusses the case of $k\geq 2,l=1$ and asks the following question:
@@
loefflerqnspecialQuestion 1  (Loeffler, see [14, Remark 4.4.2]) Assume that $k\geq 2$, $l=1$, $f$ does not have complex multiplication, and $\varepsilon_{f}\varepsilon_{g}\neq 1$. Then does there exist, for all but finitely many $\mathfrak{p}$, $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one?
Question Introduction arises in the context of Euler systems, a theory that aims to bound Selmer groups of $\mathfrak{p}$-adic Galois representations using certain classes in the Galois cohomology of its dual: a typical result is [15, Theorem 5.2.2]. For now, it seems that applying this method to a $\mathfrak{p}$-adic representation $T$ of $G_{{\mathbb{Q}}}$ requires the existence of $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T/(\sigma-1)T$ is free of rank one over the coefficient ring. For instance, in loc.cit., this is assumption (H.2). For the Galois representations considered in this article, the Euler system would typically be the Euler system of Beilinson–Flach elements constructed in [10].
Loeffler proves in [14, Proposition 4.4.1] that the answer to Question Introduction is positive when $N$ and $M$ are coprime. We prove in this article that the answer to Question Introduction is positive if certain weaker conditions are satisfied. It should be stressed that these assumptions are not verified in general, but they are reasonably easy to check in practice and apply quite broadly.
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with $q$-expansion $f=\sum_{n\geq 1}{a_{n}(f)q^{n}}$: it is well-known that there is a number field $L\subset{\mathbb{C}}$ containing $a_{n}(f)$ for every $n\geq 1$. Let $\chi$ be a primitive Dirichlet character. We say that $f$ has an inner twist by $\chi$ if there is a field embedding $\sigma:L\rightarrow{\mathbb{C}}$ and a positive multiple $N^{\prime}$ of $N$ such that, for any integer $n\geq 1$ coprime to $N^{\prime}$, $a_{n}(f)\chi(n)=\sigma(a_{n}(f))$. It can be shown (see Lemma 1) that in such a case, $\sigma(L)=L$, we can take $N^{\prime}=N$ and the conductor of $\chi$ divides $N$.
@@
If $L(E,\rho,1)\neq 0$, then the group $\mathrm{Hom}_{\widehat{\mathcal{O}_{\mathfrak{p}}}[\mathrm{Gal(K/{\mathbb{Q}})}]}(\rho,\mathrm{Sel}_{p^{\infty}}(E/K)\otimes_{{\mathbb{Z}}_{p}}\widehat{\mathcal{O}_{\mathfrak{p}}})$ is finite.
@@
    $k\geq 3$ is odd, and $f$ has an inner twist by two distinct non-trivial Dirichlet characters $\chi_{1},\chi_{2}$ such that the restrictions to $\ker{\varepsilon_{f}}$ of $\chi_{1}$ and $\chi_{2}$ are distinct.
@@
    for every distinct $g,g^{\prime}\in\mathcal{G}$, $g$ and $g^{\prime}$ are not quadratic twists of each other,
@@
    As explained in [10, Remark 11.7.5(i)], applying Theorem Introduction at a single prime shows that $\mathrm{Hom}_{\mathcal{O}[G_{{\mathbb{Q}}}]}(\rho,E(K)\otimes\mathcal{O})$ is zero. The point of Theorem Introduction is to prove that, in addition, the $\rho$-part of $\Sha(E/K)[p^{\infty}]$ is finite for certain primes $p$. In fact, after removing finitely many primes by Serre’s open image theorem [23, Théorème 3], Theorem Introduction applies to any prime $p$ which is ordinary for $E$ and such that $\mathrm{Frob}_{p}$ does not lie in the kernel of the projective representation $\tilde{\rho}$ attached to $\rho$. The first condition excludes a set of primes of density zero by [26, Cor. 2 au Théorème 20] (which is nonetheless infinite by [5, Theorem 1]), and the second condition excludes, by Chebotarev’s theorem, a set of primes of density $\frac{1}{|\mathrm{im}(\tilde{\rho})|}$. In other words, Theorem B applies to a set of primes of density $1-\frac{1}{|\mathrm{im}(\tilde{\rho})|}$, which is always at least $\frac{3}{4}$.
@@
    When $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$, and $g\in\mathcal{S}_{l}(\Gamma_{1}(M))$ are newforms with characters $\varepsilon_{f},\varepsilon_{g}$ respectively, then the Dirichlet character $\varepsilon_{f}\varepsilon_{g}$ has the same parity as $k+l$. In particular, when $k$ is even and $l=1$, $\varepsilon_{f}\varepsilon_{g}$ is odd, so it is non-trivial. The counter-examples in Theorem Introduction and Introduction are thus non trivial.
@@ 1 The big image theorem for one modular form
The results on twists are mostly taken from [12], the lemmas on inner twists come from [17], and the presentation of the results on Galois images of this section comes almost entirely from [14, Section 2].
First, we recall the definition and well-known properties of twists of newforms by Dirichlet characters. Although the author is not aware of a complete proof of the result below exactly as we state it, it is an easy consequence of the classical results of [12, §3] (on twists of modular forms by Dirichlet characters) and [16, §4.6] (about the properties of newforms).
@@
    level-twist-basicspecial $L$ divides the least common multiple $P$ of $N,CM,C^{2}$.
  3. ((((c))))
@@
    level-twist-localspecial Let $q$ be a prime number, and write $\chi=\chi_{q}\chi^{\prime}$, where the conductor of $\chi_{q}$ is a power of $q$ and that of $\chi^{\prime}$ is coprime to $q$. Then $v_{q}(L)$ only depends on $f$ and $\chi_{q}$.
@@
Proposition 1.2  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ and write its $q$-expansion $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$. Let $\sigma$ be any automorphism of ${\mathbb{C}}$. Then $\sigma(f)(\tau)=\sum_{n\geq 1}{\sigma(a_{n}(f))q^{n}}\in\mathcal{S}_{k}(\Gamma_{1}(N))$. Moreover:
  1. ((((a))))
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with primitive character $\varepsilon_{f}$. Write its $q$-expansion as $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$, and let $L={\mathbb{Q}}(a_{n}(f)\mid n\geq 1)$.
@@
Proof. –  Apart from the bound on the conductor of $\chi$, this is [17, Lemma 1.5 (i)].
Let $C$ be the conductor of $\chi$, and let $p$ be a prime such that $v_{p}(C)>v_{p}(N)$. Then by [12, Theorem 4.1] and Proposition 1((((d)))), the $p$-adic valuation of the conductor of $f\otimes\chi$ is $v_{p}(N)$, and is also $2v_{p}(C)$, and we get a contradiction. $\Box$
@@
    $\pi$ is injective with abelian image. Then the fixed field of $\mathrm{im}(\pi)\simeq\Gamma_{f}$ (which we identify to $\Gamma_{f}$) is a field $F$ such that $L/F$ is abelian with Galois group $\mathrm{im}(\pi)\simeq\Gamma_{f}$.
@@
If $\pi$ is injective, its image is an abelian subgroup of $\operatorname{Aut}(L)$. It is clear that $L/F$ is finite separable. Since $\operatorname{Aut}_{F}(L)\supset\mathrm{im}(\pi)$, one has $[L:F]\geq|\mathrm{im}(\pi)|$. Conversely, if $\alpha\in L$ is a primitive element over $F$, the monic polynomial $R(X)=\prod_{\gamma\in\mathrm{im}(\pi)}{(X-\gamma(\alpha))}\in L[X]$ vanishes at $\alpha$ and is stable under $\mathrm{im}(\pi)$, so lies in $F[X]$. Thus $[L:F]\leq[F(\alpha):F]\leq\deg{R}\leq|\mathrm{im}(\pi)|$, which concludes.
Suppose that $\pi$ is non-injective. Then there exists a non-trivial Dirichlet character $\chi$ such that $f\otimes\chi=f$. Thus $\varepsilon_{f}\chi^{2}=\varepsilon_{f}$, so $\chi$ is quadratic. The rest follows from [20, (4.4, 4.5)] in weight $k\geq 2$. $\Box$
@@
    If $c$ denotes the complex conjugation, one always has $(c,\varepsilon_{f}^{-1})\in\Gamma_{f}$ (see [20, p. 21]). As a consequence, the Galois character $\varepsilon_{f}:G_{{\mathbb{Q}}}\rightarrow L^{\times}$ factors through $\mathrm{Gal}(H/{\mathbb{Q}})$. Furthermore, the number field $F$ is fixed by the complex conjugation, so, since $L$ is totally real or CM by [20, Proposition 3.2], $F$ is totally real.
@@
Let $\mathfrak{p}\subset\mathcal{O}_{L}$ be a prime ideal with finite residue characteristic $p$. It is well-known (see [20, Theorems 2.1, 2.3]) that there exists a group homomorphism $\rho_{f,\mathfrak{p}}:G_{{\mathbb{Q}}}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$ satisfying the following properties:
@@
Remark  The $\rho_{f,\mathfrak{p}}$ that we consider is the one considered throughout the paper [20]. It is, however, dual to the one considered in [14, Definition 2.1.1]. As pointed out in [14, Remark 2.1.2], this is not important for our purposes since we are only considering the image of the representation.
Proposition 1.5  (Momose [17, Theorem 4.2], Ribet [21, §3], Ghate–González-Jiménez–Quer [7, Corollary 4.7])
Assume moreover that $k\geq 2$, that $p\nmid 2N$ is unramified in $L$ and $f$ does not have CM. Then $\rho_{f,\mathfrak{p}}(G_{H})$ has a $\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$-conjugate contained in $\{M\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}}),\,\det{M}\in{\mathbb{Z}}_{p}^{\times}\}$.
@@
Proof. –  This is [14, Theorem 2.2.2]. $\Box$
Remark  While we will not be considering it in the rest of the paper, let us briefly discuss the situation when $k\geq 2$ and $f$ has CM by a quadratic imaginary field $K$. The following claims can essentially be deduced from the contents of [20, §4] (using Serre’s theorem on rational abelian semi-simple representations [22, Chap. III., §3.1, Theorem and Remark 1]), and in particular the fact that $f$ is constructed from a Grössencharakter $\psi$ as in pp. 34–35 of op.cit..
Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}_{L}$ with residue characteristic $p$. Let $m\geq 1$ be the norm of the modulus of definition of $\psi$, and $H$ be the ray class field of $K$ modulo $m$. Assume for simplicity that $p$ is prime to $6m$ and unramified in $K/{\mathbb{Q}}$: then $H/{\mathbb{Q}}$ is unramified above $p$. Then $(\rho_{f,\mathfrak{p}})_{|G_{K}}$ is abelian, one has $\rho_{f,\mathfrak{p}}(I_{K,p})=\rho_{f,\mathfrak{p}}(G_{H})$, where $I_{K,p}$ is the subgroup of $G_{K}^{ab}$ generated by the inertia subgroups of prime ideals of $K$ above $p$.
Moreover, let $\mathfrak{a}:\mathbb{A}_{K}^{\times}/K^{\times}\rightarrow G_{K}^{ab}$ denote the class field theory reciprocity homomorphism (mapping a uniformizer to an arithmetic Frobenius), then one has $\mathfrak{a}((\mathcal{O}_{K}\otimes{\mathbb{Z}}_{p})^{\times})=I_{K,p}$ and $\left(\rho_{f,\mathfrak{p}}\circ\mathfrak{a}\right)_{|(\mathcal{O}_{K}\otimes{\mathbb{Z}}_{p})^{\times}}$ is conjugate to the $(1-k)$-th power of the representation

### v3

@@ Introduction
Let $k,N\geq 1$ be integers. We denote by $\mathcal{S}_{k}(\Gamma_{1}(N))$ the space of cuspidal holomorphic modular forms with level $\Gamma_{1}(N)$. Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a cuspidal newform with character $\varepsilon_{f}$, and $L$ be a number field containing its coefficients. Given a maximal ideal $\mathfrak{p}\subset\mathcal{O}_{L}$, we can attach to $f$ a free $\mathcal{O}_{L_{\mathfrak{p}}}$-module $T_{f,\mathfrak{p}}$ of rank two endowed with a continuous linear action of $G_{{\mathbb{Q}}}:=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})$. By works of Momose [18] and Ribet [22], it is known that, when $f$ has no complex multiplication and $k\geq 2$, the image of $G_{{\mathbb{Q}}}$ in $\operatorname{Aut}(T_{f,\mathfrak{p}})\simeq\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$ contains a conjugate of $\operatorname{SL}_{2}({\mathbb{Z}}_{p})$ for all but finitely many $\mathfrak{p}$. More recently, Loeffler [15] proved an adelic open image theorem for such newforms analogous to that of Serre [24] for elliptic curves without complex multiplication.
@@
then, for all but finitely many $\mathfrak{p}$, there exists a $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}:=\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}}(\mu_{p^{\infty}}))$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one. Loeffler also discusses the case where $k\geq 2,l=1$ and asks the following question:
@@
loefflerqnspecialQuestion 1  (Loeffler, see [15, Remark 4.4.2]) Assume that $k\geq 2$, $l=1$, $f$ does not have complex multiplication, and $\varepsilon_{f}\varepsilon_{g}\neq 1$. Is it true that, for all but finitely many $\mathfrak{p}$, there exists $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T_{f,g,\mathfrak{p}}/(\sigma-1)T_{f,g,\mathfrak{p}}$ is free of rank one?
Question Introduction arises in the context of Euler systems, a theory that aims to bound Selmer groups of $\mathfrak{p}$-adic Galois representations using certain classes in the Galois cohomology of its dual: a typical result is [16, Theorem 5.2.2]. For now, it seems that applying this method to a $\mathfrak{p}$-adic representation $T$ of $G_{{\mathbb{Q}}}$ requires the existence of $\sigma\in G_{{\mathbb{Q}}(\mu_{p^{\infty}})}$ such that $T/(\sigma-1)T$ is free of rank one over the coefficient ring. For instance, in loc.cit., this is assumption (H.2). For the Galois representations considered in this article, the Euler system would typically be the Euler system of Beilinson–Flach elements constructed in [10].
Loeffler proves in [15, Proposition 4.4.1] that the answer to Question Introduction is positive when $N$ and $M$ are coprime. We prove in this article that the answer to Question Introduction is positive if certain weaker conditions are satisfied. It should be stressed that these assumptions are not verified in general, but they are reasonably easy to check in practice and apply quite broadly.
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with $q$-expansion $f=\sum_{n\geq 1}{a_{n}(f)q^{n}}$: it is well-known that there is a number field $L\subset{\mathbb{C}}$ containing $a_{n}(f)$ for every $n\geq 1$. Let $\chi$ be a primitive Dirichlet character. We say that $f$ has an inner twist by $\chi$ if there is a field embedding $\sigma:L\rightarrow{\mathbb{C}}$ and a positive multiple $N^{\prime}$ of $N$ such that, for any integer $n\geq 1$ coprime to $N^{\prime}$, $a_{n}(f)\chi(n)=\sigma(a_{n}(f))$. It can be shown (see Lemma 1) that, in such a case, the conductor of $\chi$ divides $N$, we can take $N^{\prime}=N$, and one has $\sigma(L)=L$.
@@
[…]
@@
    $k\geq 3$ is odd, and $f$ has an inner twist by two distinct non-trivial Dirichlet characters $\chi_{1},\chi_{2}$ such that the restrictions to $\ker{\varepsilon_{f}}$ of $\chi_{1}$ and $\chi_{2}$ are both non-trivial and distinct.
@@
    for every distinct $g^{\prime}\in\mathcal{G}, gg$ and $g^{\prime}$ are not quadratic twists of each other,
@@
    As explained in [10, Remark 11.7.5(i)], applying Theorem Introduction at a single prime shows that $\mathrm{Hom}_{\mathcal{O}[G_{{\mathbb{Q}}}]}(\rho,E(K)\otimes\mathcal{O})$ is zero. The point of Theorem Introduction is to prove that, in addition, the $\rho$-part of $\Sha(E/K)[p^{\infty}]$ is finite for certain primes $p$. In fact, after removing finitely many primes by Serre’s open image theorem [24, Théorème 3], Theorem Introduction applies to any prime $p$ which is ordinary for $E$ and such that $\mathrm{Frob}_{p}$ does not lie in the kernel of the projective representation $\tilde{\rho}$ attached to $\rho$. The first condition excludes a set of primes of density zero by [27, Cor. 2 au Théorème 20] (which is nonetheless infinite by [5, Theorem 1]), and the second condition excludes, by Chebotarev’s theorem, a set of primes of density $\frac{1}{|\mathrm{im}(\tilde{\rho})|}$. In other words, Theorem B applies to a set of primes of density $1-\frac{1}{|\mathrm{im}(\tilde{\rho})|}$, which is always at least $\frac{3}{4}$.
@@
    When $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$, and $g\in\mathcal{S}_{l}(\Gamma_{1}(M))$ are newforms with characters $\varepsilon_{f},\varepsilon_{g}$ respectively, the Dirichlet character $\varepsilon_{f}\varepsilon_{g}$ has the same parity as $k+l$. In particular, when $k$ is even and $l=1$, $\varepsilon_{f}\varepsilon_{g}$ is odd, so it is non-trivial. Thus, the counter-examples in Theorems Introduction and Introduction are not trivial.
@@ 1 The big image theorem for one modular form
The results on twists are mostly taken from [13], the lemmas on inner twists come from [18], and the presentation of the results on Galois images of this section comes almost entirely from [15, Section 2].
We start with twists of newforms by Dirichlet characters. Although the author is not aware of a complete proof of the result below exactly as we state it, it is an easy consequence of the classical results of [13, §3] (on twists of modular forms by Dirichlet characters) and [17, §4.6] (about the properties of newforms).
@@
    level-twist-basicspecial $L$ divides the least common multiple of $N,CM,C^{2}$.
  3. ((c))
@@
    level-twist-localspecial Let $q$ be a prime number, and write $\chi=\chi_{q}\chi^{\prime}$, where the conductor of $\chi_{q}$ (resp. $\chi^{\prime}$) is a power of $q$ (resp. coprime to $q$). Then $v_{q}(L)$ only depends on $f$ and $\chi_{q}$.
@@
Proposition 1.2  Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ and write its $q$-expansion $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$. Let $\sigma$ be any automorphism of $\mathbb{C}$}. Then $\sigma(f)(\tau):=\sum_{n\geq 1}{\sigma(a_{n}(f))q^{n}}\in\mathcal{S}_{k}(\Gamma_{1}(N))$. Moreover:
  1. ((a))
@@
Let $f\in\mathcal{S}_{k}(\Gamma_{1}(N))$ be a newform with (primitive) character $\varepsilon_{f}$. Let $f(\tau)=\sum_{n\geq 1}{a_{n}(f)q^{n}}$ be its $q$-expansion and let $L={\mathbb{Q}}(a_{n}(f)\mid n\geq 1)$.
@@
Proof. –  Apart from the bound on the conductor of $\chi$, this is [18, Lemma 1.5 (i)].
Let $C$ be the conductor of $\chi$, and let $p$ be a prime such that $v_{p}(C)>v_{p}(N)$. Then by [13, Theorem 4.1] and Proposition 1((d)), the $p$-adic valuation of the conductor of $f\otimes\chi$ is $v_{p}(N)$, and is also $2v_{p}(C)$, and we get a contradiction. $\Box$
@@
    $\pi$ is injective with abelian image. Let $F$ be the subfield of $L$ fixed by $\mathrm{im}(\pi)$ (which we identify to $\Gamma_{f}$): the extension $L/F$ is abelian with Galois group $\mathrm{im}(\pi)\simeq\Gamma_{f}$.
@@
If $\pi$ is injective, its image is an abelian subgroup of $\operatorname{Aut}(L)$. Let $F$ be the field fixed by $\mathrm{im}(\pi)$, then it is well-known (e.g. [11, VI, Theorem 1.8]) that $L/F$ is Galois with Galois group $\mathrm{im}(\pi)\simeq\Gamma_{f}$ which is abelian.
Suppose that $\pi$ is not injective. Then there exists a non-trivial Dirichlet character $\chi$ such that $f\otimes\chi=f$. Thus $\varepsilon_{f}\chi^{2}=\varepsilon_{f}$, so $\chi$ is quadratic. The rest follows from [21, (4.4, 4.5)] in weight $k\geq 2$. $\Box$
@@
    If $c$ denotes the complex conjugation, one always has $(c,\varepsilon_{f}^{-1})\in\Gamma_{f}$ (see [21, p. 21]). As a consequence, the Galois character $\varepsilon_{f}:G_{{\mathbb{Q}}}\rightarrow L^{\times}$ factors through $\mathrm{Gal}(H/{\mathbb{Q}})$. Furthermore, the number field $F$ is fixed by the complex conjugation, so, since $L$ is totally real or CM by [21, Proposition 3.2], $F$ is totally real.
@@
Let $\mathfrak{p}\subset\mathcal{O}_{L}$ be a prime ideal with finite residue characteristic $p$. It is well-known (see [21, Theorems 2.1, 2.3]) that there exists a group homomorphism $\rho_{f,\mathfrak{p}}:G_{{\mathbb{Q}}}\rightarrow\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$ satisfying the following properties:
@@
Remark  The $\rho_{f,\mathfrak{p}}$ that we consider is the one considered throughout the paper [21]. It is, however, dual to the one considered in [15, Definition 2.1.1]. As pointed out in [15, Remark 2.1.2], this is not important for our purposes since we are only considering the image of the representation.
Proposition 1.5  (Momose [18, Theorem 4.2], Ribet [22, §3], Ghate–González-Jiménez–Quer [7, Corollary 4.7])
Assume that $k\geq 2$, that $p\nmid 2N$ is unramified in $L$ and $f$ does not have CM. Then $\rho_{f,\mathfrak{p}}(G_{H})$ has a $\operatorname{GL}_{2}(\mathcal{O}_{L_{\mathfrak{p}}})$-conjugate contained in $\{M\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}}),\,\det{M}\in{\mathbb{Z}}_{p}^{\times}\}$.
@@
Proof. –  This is [15, Theorem 2.2.2]. $\Box$
Remark  While we will not be considering it in the rest of the paper, let us briefly discuss the situation when $k\geq 2$ and $f$ has CM by an imaginary quadratic field $K$. The following claims can essentially be deduced from the contents of [21, §4] (using Serre’s theorem on rational abelian semi-simple representations [23, Chap. III., §3.1, Theorem and Remark 1]), and in particular the fact that $f$ is constructed from a Grössencharakter $\psi$ as in [21, pp. 34–35].
Let $\mathfrak{p}$ be a maximal ideal of $\mathcal{O}_{L}$ with residue characteristic $p$. Let $m\geq 1$ be the norm of the modulus of definition of $\psi$, and $H$ be the ray class field of $K$ modulo $m$. Assume for simplicity that $p$ is prime to $6m$ and unramified in $K/{\mathbb{Q}}$. Then $H/{\mathbb{Q}}$ is unramified above $p$, $(\rho_{f,\mathfrak{p}})_{|G_{K}}$ is abelian, and one has $\rho_{f,\mathfrak{p}}(I_{K,p})=\rho_{f,\mathfrak{p}}(G_{H})$, where $I_{K,p}$ is the subgroup of $G_{K}^{ab}$ generated by the inertia subgroups of prime ideals of $K$ above $p$.
Moreover, let $\mathfrak{a}:\mathbb{A}_{K}^{\times}/K^{\times}\rightarrow G_{K}^{ab}$ denote the class field theory reciprocity homomorphism (mapping a uniformizer to an arithmetic Frobenius), then one has $\mathfrak{a}((\mathcal{O}_{K}\otimes{\mathbb{Z}}_{p})^{\times})=I_{K,p}$ and $\left(\rho_{f,\mathfrak{p}}\circ\mathfrak{a}\right)_{|(\mathcal{O}_{K}\otimes{\mathbb{Z}}_{p})^{\times}}$ is conjugate to the (1k)(1-k)-th power of the representation